OpenStudy (anonymous):

Which equation has the solution 4x - 12 = 32

OpenStudy (uri):

4x - 12 = 32 Add 12 both sides 4x - 12 = 32 +12 +12 ========== 4x=44 x=44/4 x=11

jhonyy9 (jhonyy9):

11 is ok